1. 2X + 5 = 17

2. 2X + 3X - 5 = 25

Answer by Abbey(339)

2x+5=17
subtract 5 from both sides:
2x=12
divide both sides by 2:
x=6
2x+3x-5=25
Combine like terms:
5x-5=25
Add 5 to both sides
5x=30
divide both sides by 5
x=6